Gaussian divergence theorem

The velocity field of a fluid is given by F (?, ?, ?) = 5?k and let ? be the sphere ?2 + ?2 + ?2 = 16. Find the flow from ? through ?.
Answer #1

Use the divergence theorem:
p = 5 = density
Mass flux = p ???r div(v) dVol
=p ???r div(-yi + xj + 2zk) dVol
= p ???r (-1 + 1+ 2 ) dVol
= 2p ?1 dVol
r^2 = x^2 + y^2 + z^2 = 16
r = 4 units
==>Vol = [0,r=4] ? 1 dVol = (4/3)?43
Mass flux = 2*p*(4/3)?43
Mass flux = 853.33? units/second

answered by: ANURANJAN SARSAM